Reymart is 5 times as old as Jonel in 9 years, Reymart will twice as old as Jonel. What are their ages now?
Answer by venugopalramana(3286)

Reymart (R say)is 5 times as old as Jonel (J SAY)
in 9 years, Reymart will twice as old as Jonel. What are their ages now?
LET AGE OF R BE R YEARS AND AGE OF J BE J YEARS NOW
R/J=5
R=5J.............................1
IN 9 YEARS
THEIR AGES WILL BE R+9 AND J+9 YEARS.HENCE
(R+9)/(J+9)=2
R+9=2(J+9)=2J+18.............................2
SUBSITUTING FROM EQN.1 FOR R
5J+9=2J+18
5J-2J=3J=18-9=9
J=3
R=5*3=15